# Approximation in Lattice Field Theories

### v1

@@ 4 Results
Figure 7: Plot of $\langle x(1)^{2}\rangle$ estimates for $N=10^{3}$ and $M=9$ with error bars. The number of sweeps before adding a measurement to the Markov Chain $h_{*}$ was around 66 for the $\langle x(i)^{2}\rangle$ functions, including this one. Thus, the Markov Chain for this correlator was around 33,000 elements in length.
The methodology presented in Section 3 is viable for higher values of $M$, but there may be some nuance to consider. Firstly, higher values of $M$ (ones larger than our examples) correspond to lattices with many more sites. Thus, $M\gg 1$ leads to the classical case. This level of discretization may make the autocorrelations fall much slower, resulting in the need for a slightly modified algorithm using microcanonical updating. Rather than updating sites with some $\delta$, the next value at the site is determined by the properties of the action in this updating technique. More information on its theory and implementation can be found in [4]. Other ways to reduce error can also be found in [3]. For our purposes, the number of observables grows as $M^{2}$, so considering $M\geq 16$ is computationally costly. This is especially true since we are running the code with $2\times 10^{5}$ measurements - a value that is high enough to minimize error while being small enough to run for $M=4$ and $M=9$. More specifically, the error for all correlators is of order $10^{-3}$, since:
@@
Above, we used the fact that the variance of the Markov Chain $X$ is approximately the autocorrelation threshold we set, which was $0.1$. We noticed that correlation functions of the form $\langle x^{2}(i)\rangle$ had larger spacing $h_{*}$ leading to larger errors of order $10^{-3}$ or smaller errors of order $10^{-2}$.
The approximation can be verified in two separate ways. Firstly, we can simply plug in the 1 and 2-pt connected correlation values into the right hand side of eqns. (4) and 5. The closer these are to 0, then the better the approximation was. The approximation is thus successful if these equations are well-satisfied for large enough $N$. The mathematica file that we used for this purpose for $M=9$ can be found on GitHub. A more analytical approach is to consider only the $O(1)$ terms of the right hand side of the aforementioned equations. Then, deriving values for $W_{i}$ and $W_{ij}$ is straightforward, and we can compare these with the numerically calculated values. Both of these methods are fundamentally the same by the nature of the problem, but we will utilize both methods to avoid any loose ends. The trend we can also look for is the right hand side values of eqns. (4) and 5 must get smaller for larger $N$. As $N$ grows larger, the numerically obtained values should draw closer to the derived values for the approximation to succeed. However, we expect the error of all of our values to be of order $10^{-3}$. Thus, we must keep in mind that if the equations are well-satisfied to order $10^{-3}$ or below, then the proposed approximation works well. Similarly, a discrepancy up to order $10^{-3}$ between the exact values derived and the numerically evaluated correlators is reasonable. This precision can be fine-tuned simply by the number of measurements.
@@
$N=10^{2}$ $\mathbb{1}$ $9.13\times 10^{-3}$ $1.11\times 10^{-2}$ $2.00\times 10^{-2}$ $2.11\times 10^{-2}$
@@
$N=10^{3}$ $\mathbb{1}$ $1.06\times 10^{-2}$ $1.53\times 10^{-2}$ $1.09\times 10^{-2}$ $5.37\times 10^{-3}$
@@
$N=10^{6}$ $\mathbb{1}$ $3.38\times 10^{-3}$ $2.31\times 10^{-3}$ $5.01\times 10^{-3}$ $2.53\times 10^{-3}$
$x(1)$ $1.063$ $1.35\times 10^{-1}$ $1.38\times 10^{-1}$ $1.90\times 10^{-1}$
$x(2)$ - $1.067$ $1.96\times 10^{-1}$ $1.33\times 10^{-1}$
$x(3)$ - - $1.062$ $1.38\times 10^{-1}$
$x(4)$ - - - $1.060$
Table 1: Table of 1 and 2-pt correlation functions for $N=10,10^{3}$, and $10^{6}$, $M=4$, and $\Delta=2.75$, where the element in row $x(i)$ and column $x(j)$ corresponds to $\langle x(i)x(j)\rangle$. Repeat values are not included for visibility. The error for all of these values were of order $10^{-3}$ , except the error of correlation functions $\langle x(i)^{2}\rangle$. These error larger by magnitude $\sim\sqrt{2}$ due to autocorrelation spacing, so their error was close to order $10^{-2}$. Note that the 1-pt correlation functions approach 0 for larger $N$. The same can be said about the 2-pt functions, though the diagonal elements have higher error, so it is not as obvious.
@@ 4.1 $M=4$
The specific values of the correlation functions for $M=4$ can be found in Table 1. A straightforward calculation using the $\sqrt{N}$ term of the action revealed $\Phi_{0}=-0.0709$. However, $\det{(K_{ij}+(\Phi_{0}+\delta)\mathbb{1})}$ is negative [^4] Here, $K_{ij}$ is a 2D discrete Laplacian plus an identity matrix, as described in Section 3.2., where $\delta$ here is any form of fluctuation. This poses a problem as the action becomes complex. Since the region around the stationary point is the only region necessary to evaluate our desired correlators, we can instead consider the absolute value (or equivalently, changing sign, which is what we did in the code). We are simply ridding the action of a complex constant $\pm i\pi$, which is negligible.
Table 2 contains what the $4$ and $10$ expressions respectively evaluated to, the maximum of which acts as a metric for this approximation. One can easily notice the success of the approximation at $N\sim 10^{3}$ (or even $N\sim 100$, but we are trying to be conservative) and beyond. In addition, the values of Table 2 rightfully decrease as $N$ grows.
In addition, we can determine if the numerically found values approach derived $W_{i}$ and $W_{ij}$ for large $N$. In eqn. (4), only the first term is relevant. This term vanishes if and only if all 1-pt correlation functions vanish. Table 1 shows this behavior; as $N$ grows larger, $\langle x(i)\rangle$ approach 0. The other set of equations from eqn. (5) require the 2-pt correlation functions to have the following values:
$\langle x(i)x(j)\rangle=W_{ij}=\begin{pmatrix}1.06152&0.138862&0.138862&0.193632\\ -&1.06152&0.193632&0.138862\\ -&-&1.06152&0.138862\\ -&-&-&1.06152\\ \end{pmatrix}$. (16)
Table 1 confirms this for the off-diagonal terms on average. A pattern for the diagonal terms may be less obvious, and this discrepancy is most likely due to the number of measurements used in their Markov Chains. The numerical analysis suggested that around two times as many sweeps were needed to reduce the autocorrelation to less than $0.1$. Thus, the error for these terms were near $10^{-2}$. These results strongly suggest that the approximation holds true for a $2\times 2$ lattice from $N\sim 10^{3}$ onward.
@@
$N=10^{2}$ N/A $1.04\times 10^{-2}$ $1.26\times 10^{-2}$ $2.28\times 10^{-2}$ $2.42\times 10^{-2}$
11 $2.21\times 10^{-3}$ $4.27\times 10^{-3}$ $1.51\times 10^{-3}$ $1.64\times 10^{-3}$
22 - $7.79\times 10^{-3}$ $3.49\times 10^{-3}$ $5.03\times 10^{-3}$
33 - - $3.27\times 10^{-3}$ $2.64\times 10^{-3}$
44 - - - $4.80\times 10^{-3}$
$N=10^{3}$ N/A $9.46\times 10^{-3}$ $1.42\times 10^{-2}$ $9.14\times 10^{-3}$ $3.45\times 10^{-3}$
11 $1.93\times 10^{-3}$ $4.59\times 10^{-3}$ $2.78\times 10^{-5}$ $4.74\times 10^{-3}$
22 - $9.64\times 10^{-3}$ $3.96\times 10^{-3}$ $2.89\times 10^{-3}$
33 - - $7.53\times 10^{-3}$ $1.59\times 10^{-3}$
44 - - - $5.63\times 10^{-4}$
$N=10^{6}$ N/A $3.71\times 10^{-3}$ $1.99\times 10^{-3}$ $5.10\times 10^{-3}$ $2.74\times 10^{-3}$
11 $2.34\times 10^{-3}$ $3.25\times 10^{-3}$ $9.47\times 10^{-4}$ $2.49\times 10^{-3}$
22 - $5.93\times 10^{-3}$ $2.28\times 10^{-3}$ $5.52\times 10^{-3}$
33 - - $2.03\times 10^{-4}$ $7.06\times 10^{-4}$
44 - - - $4.15\times 10^{-4}$
Table 2: Table of right-hand side expressions of eqns. (4) and (5) with the calculated 1 and 2-pt correlation functions for $N=10,10^{3},$ and $10^{6}$. The first row of each NN value corresponds to the four 1st order equations, while the last four rows correspond to the 10 unique 2nd order equations. Note that the closer they are to 0, the more the aforementioned equations are true - validating the approximation. It is clear to see that the maximum value for NN decreases as NN increases, which strongly suggests that the approximation is true for large $N$.
@@ 4.2 $M=9$
For the $3\times 3$ lattice, there are nine 1 and forty-five 2-pt correlation functions. Like in the previous section, we considered $N=10^{2},10^{3},$and $10^{6}$. The stationary point for this system was found to be $\Phi_{0}=0.261$. Their values, along with what the expressions of the approximation evaluate to, can be found in Table A1. It is clear that the equations are much more satisfied at $N=10^{3}$ than $N=10^{2}$. The difference when one compares $N=10^{3}$ and $10^{6}$ is less obvious. If the maximum is our metric like in the $M=4$ case, then $N=10^{6}$ is only a slight improvement. However, we must not disregard the higher impact of error in this setting. The important thing is the following: we see that all the equations are satisfied for $N=10^{3}$ and above up to order $10^{-3}$. This level of accuracy is what we hoped for earlier. Thus, these values strongly suggest that the approximation holds true on a $3\times 3$ lattice for $N\sim 10^{3}$ (again, we are trying to be conservative) and higher.
We can also derive $W_{i}$ and $W_{ij}$ in large $N$ limit to see if this gives any extra perspective. As expected, the 1-pt correlators must vanish for large $N$. The exact values for the 2-pt correlators are:
$W_{ij}=10^{-2}\begin{pmatrix}108.1&1.098&1.098&1.098&4.63&4.63&1.098&4.63&4.63\\ -&108.1&1.098&4.63&1.098&4.63&4.63&1.098&4.63\\ -&-&108.1&4.63&4.63&1.098&4.63&4.63&1.098\\ -&-&-&108.1&1.098&1.098&1.098&4.63&4.63\\ -&-&-&-&108.1&1.098&4.63&1.098&4.63\\ -&-&-&-&-&108.1&4.63&4.63&1.098\\ -&-&-&-&-&-&108.1&1.098&1.098\\ -&-&-&-&-&-&-&108.1&1.098\\ -&-&-&-&-&-&-&-&108.1\\ \end{pmatrix}$. (17)
Like the equations, the numerical 1 and 2-pt correlators for $N=10^{3}$ match the above values better than for $N=10^{2}$. However, we can notice that some correlation functions of the form $\langle x^{2}(i)\rangle$ for $N=10^{6}$ grow further away from the above values compared to $N=10^{3}$. This is most likely due to error; we found that these type of functions had larger spacing $h_{*}$, leading to more error, just like in the previous lattice sizing. Other than this, the $N=10^{6}$ values are closer on average to the ones above.
@@ 5 Conclusion
We’ve seen that the proposed level of SD truncation gives excellent results on very small square lattices. This encourages us to proceed with the main physics objective that motivated the invention of these approximation schemes. The homogeneous electron fluid is the basic model on which condensed matter physics is built. Density Functional Theory (DFT) is an attempt to write down phenomenological ansatze for the correlation functions of this field theory, in order to understand real materials. From the earliest days of study of the HEF it was realized that it had both a fluid and a crystalline phase. More recently[5] colloidal phases have been proposed, and a persuasive mean field argument for the existence of these phases has been given in two dimensions.
In a series of papers [6, 7, 8], one of the authors has tried to explore these colloidal phases more systematically. In [2], he proposed that one could do so by searching for crystalline solutions of the truncated large $N$ SD equations. These would provide a more systematic approach to the Wigner crystal phase than the complicated semi-classical solution of the many body Schrodinger equation and allow an extrapolation of the solution into the region where the crystal became meta-stable. As argued in [6, 7, 8] this is the regime where one expects gel like colloids to exist. One should also be able to follow the system through its metal-insulator transition and check whether the Kivelson-Spivak prediction of crystalline impurities with negative surface tension is valid at large enough $N$. Finally, the disappearance of the crystalline solution at a finite value of the density would be a neat explanation of the mysterious phase transition unveiled in [9, 10].

### v2

@@ 4 Results
Figure 7: Plot of $\langle x(1)^{2}\rangle$ estimates for $N=1000$ and $M=9$ with error bars. The number of sweeps before adding a measurement to the Markov Chain $h_{*}$ was around 66 for the $\langle x(i)^{2}\rangle$ functions, including this one. Thus, the Markov Chain for this correlator was around 33,000 elements in length.
The methodology presented in section 3 is viable for higher values of $M$, but there may be some nuance to consider. Firstly, higher values of $M$ (ones larger than our examples) correspond to lattices with many more sites. Thus, $M\gg 1$ leads to the classical case. This level of discretization may make the autocorrelations fall much slower, resulting in the need for a slightly modified algorithm using microcanonical updating. Rather than updating sites with some $\delta$, the next value at the site is determined by the properties of the action in this updating technique. More information on its theory and implementation can be found in morningstar2007monte. Other ways to reduce error can also be found in budd2023monte. For our purposes, the number of observables grows as $M^{2}$, so considering $M\geq 16$ is computationally costly. This is especially true since we are running the code with $2\times 10^{5}$ measurements - a value that is high enough to minimize error while being small enough to run for $M=4$ and $M=9$. More specifically, the error for all correlators is of order $10^{-3}$, since:
@@
Above, we used the fact that the variance of the Markov Chain $X$ is approximately the autocorrelation threshold we set, which was $0.1$. We noticed that correlation functions of the form $\langle x^{2}(i)\rangle$ had larger spacing $h_{*}$ leading to errors high in the $10^{-3}$’s or low in the $10^{-2}$’s.
For both cases of $M$, we choose to find correlation functions for $N=40,100,$ and $1000$. Larger $N$ can certainly be studied, but we found that $N=1000$ is around the threshold in which the equations are satisfied to order $10^{-3}$. Since the error values are also at a similar level, the equations would be more prone to error and less insightful for values of $N$ larger than $1000$. On the other hand, we chose $N=40$ as our baseline. Lower values of $N$ led to the determinant becoming 0 or negative (from eq. (1)), when we retained a probability of acceptance of $50\%-60\%$. We found that $N=40$ is large enough to safely avoid this issue.
The approximation can be verified by simply plugging in the 1 and 2-pt correlation values into the right hand sides of eqs. (4) and (5). The closer these are to 0, the better the approximation is. The Mathematica file that we used for this purpose for $M=9$ can be found on GitHub. The trend we must also look for is that these right hand side values should get smaller for larger $N$. However, we expect the error of all of our values to be of order $10^{-3}$. Thus, we must keep in mind that if the equations are well-satisfied to about order $10^{-3}$ or below, then the proposed approximation works well [^4] This precision can be fine-tuned by increasing the number of measurements.. When analyzing the results for these choices of $N$, we must keep in mind that we do not expect to see one drastically outperforming another. The $1/N$ term for $N=40$ will be $\sim 25$ times larger than for $N=1000$, and the difference between the $1/\sqrt{N}$ terms will only be off by a magnitude of $5$. This difference will be even smaller when considering the other two pairs of $N$ values. Thus, any clear decrease in the right hand sides of eqs. (4) and (5) points to a successful approximation.
A more roundabout way to check the approximation is to derive the values of the 1 and 2-pt correlators that would satisfy the leading order ($\mathcal{O}(1)$) terms of eqs. (4) and (5). Doing this reveals that $W_{i}$ should approach 0 for larger $N$. This can be utilized as a simple, visual check when comparing different values of $N$. The 2-pt correlation functions are less trivial, so we will not concern ourselves with any comparisons for these functions.
@@
$N=40$ $\mathbb{1}$ $1.35\times 10^{-2}$ $26\times 10^{-2} 1.1$$18\times 10^{-2}$ $1.76\times 10^{-2}$
$x(1)$ $1.068$ $1.49\times 10^{-1}$ $1.52\times 10^{-1}$ $2.09\times 10^{-1}$
$x(2)$ - $1.073$ $1.98\times 10^{-1}$ $1.48\times 10^{-1}$
$x(3)$ - - $1.077$ $1.40\times 10^{-1}$
$x(4)$ - - - $1.063$
$N=100$ $\mathbb{1}$ $9.13\times 10^{-3}$ $1.11\times 10^{-2}$ $2.00\times 10^{-2}$ $2.11\times 10^{-2}$
@@
$N=1000$ $\mathbb{1}$ $1.06\times 10^{-2}$ $1.53\times 10^{-2}$ $1.09\times 10^{-2}$ $5.37\times 10^{-3}$
@@
Table 1: Table of 1 and 2-pt correlation functions for $N=40,100,$ and $1000$, $M=4$, and $\Delta=2.75$, where the element in row $x(i)$ and column $x(j)$ corresponds to $\langle x(i)x(j)\rangle$. Repeat values are not included for visibility. The error for all of these values were of order $10^{-3}$ , except the error of correlation functions $\langle x(i)^{2}\rangle$. These error larger by magnitude $\sim\sqrt{2}$ due to autocorrelation spacing, so their error was close to order $10^{-2}$.
@@ 4.1 $M=4$
The specific values of the correlation functions for $M=4$ for $N=40,100$, and $1000$ can be found in table 1. A straightforward calculation using the $\sqrt{N}$ term of the action revealed $\Phi_{0}=-0.0709$. However, $\det{(K_{ij}+(\Phi_{0}+\delta)\mathbb{1})}$ is negative, where $\delta$ here is any form of fluctuation. [^5] Here, $K_{ij}$ is a 2D discrete Laplacian plus an identity matrix, as described in section 3.2. This poses a problem as the action becomes complex. Since the region around the stationary point is the only region necessary to evaluate our desired correlators, we can instead consider the absolute value (or equivalently, changing sign, which is what we did in the code). We are simply ridding the action of a complex constant $\pm i\pi$, which is negligible.
Table 2 contains what the four and ten expressions respectively evaluated to. Immediately, we can tell that the approximation works somewhat-well even at $N=40$, but the equations are better-satisfied at higher values of $N$: there are many entries of order $10^{-2}$ for $N=40$, but only one such entry exists for $N=100$ and $1000$ respectively. The increased number of terms of order $10^{-4}$ and the lower coefficients for terms of order $10^{-3}$ of $N=1000$ suggests that the approximation is better for larger $N$. Observing the 1-pt correlators in table 1 also indicates this; these values also get smaller for larger $N$. Both of these results strongly suggest that the approximation holds true for a $2\times 2$ lattice from $N\sim 100$ onward.
@@
$N=40$ N/A $4.16\times 10^{-3}$ $4.91\times 10^{-3}$ $5.96\times 10^{-3}$ $3.58\times 10^{-4}$
11 $1.25\times 10^{-2}$ $6.26\times 10^{-3}$ $2.97\times 10^{-3}$ $5.29\times 10^{-4}$
22 - $5.31\times 10^{-3}$ $1.26\times 10^{-2}$ $5.89\times 10^{-3}$
33 - - $1.42\times 10^{-3}$ $1.60\times 10^{-2}$
44 - - - $1.65\times 10^{-2}$
$N=100$ N/A $4.63\times 10^{-3}$ $2.50\times 10^{-3}$ $7.83\times 10^{-3}$ $9.12\times 10^{-3}$
11 $1.66\times 10^{-3}$ $3.89\times 10^{-5}$ $2.76\times 10^{-3}$ $5.84\times 10^{-3}$
22 - $1.17\times 10^{-2}$ $7.73\times 10^{-3}$ $9.39\times 10^{-3}$
33 - - $7.40\times 10^{-3}$ $1.92\times 10^{-3}$
44 - - - $6.06\times 10^{-4}$
$N=1000$ N/A $5.04\times 10^{-3}$ $9.80\times 10^{-3}$ $4.75\times 10^{-3}$ $9.46\times 10^{-4}$
11 $2.37\times 10^{-3}$ $5.10\times 10^{-3}$ $5.10\times 10^{-4}$ $5.18\times 10^{-3}$
22 - $1.02\times 10^{-2}$ $4.49\times 10^{-3}$ $3.36\times 10^{-3}$
33 - - $7.98\times 10^{-3}$ $1.15\times 10^{-3}$
44 - - - $1.91\times 10^{-4}$
Table 2: Table of right-hand side expressions of eqs. (4) and (5) with the calculated 1 and 2-pt correlation functions for $N=40,100,$ and $1000$. The first row of each NN value corresponds to the four 1st order equations, while the last four rows correspond to the 10 unique 2nd order equations. Note that the closer they are to 0, the more the aforementioned equations are true - validating the approximation. It is clear to see that the maximum value for NN decreases as NN increases, which strongly suggest that the approximation is true for large $N$.
@@ 4.2 $M=9$
For the $3\times 3$ lattice, there are nine $1$ and forty-five $2$-pt correlation functions. As before, we considered $N=40,100,$ and $1000$. The stationary point for this system was found to be $\Phi_{0}=0.261$. Their values, along with what the expressions of the approximation evaluate to, can be found in table LABEL:tbl:_M_=_9. At an initial glance, we see that the 1-pt correlation functions decrease in magnitude and approach 0 for larger $N$. This fact is a good indicator of our approximation’s success for this lattice.
When we look at the results from eq. (4), we see that the $N=40$ and $N=1000$ cases are satisfied well. The $N=100$ case does this to a slightly lesser extent, due to the $2$ extra entries of order $10^{-2}$ (which are right on the cusp of $10^{-3}$). Before making a conclusion, we must also consider the forty-five remaining values from eq. (5). Here, we find a consistent improvement for larger $N$. Namely, the number of terms of order $10^{-2}$ and $10^{-3}$ decrease as $3\rightarrow 3\rightarrow 2$ and $38\rightarrow 34\rightarrow 32$ respectively for $N=40,100$, and $1000$. These missing terms appear as $4\rightarrow 7\rightarrow 10$ values of order $10^{-4}$ and $0\rightarrow 1\rightarrow 1$ values of order $10^{-5}$.
While the first nine equation results may seem worrisome, we must keep in mind how close in proximity the values of $N$ are. This reason is why the 1 and 2-pt function values for $N=40$ and $N=100$ are similar, though the latter is slightly better overall. The more telling scenario is when we consider the $N=40$ and $N=1000$ values; the latter terms satisfy the equations much better. When assessing both sets of equations and their results, we see that there is an overall improvement for larger $N$. All of this analysis signifies that the approximation holds true on a $3\times 3$ lattice for $N\sim 100$ and higher.
@@ 5 Conclusion
We’ve seen that the proposed level of Schwinger-Dyson (SD) truncation gives excellent results on very small square lattices. This encourages us to proceed with the main physics objective that motivated the invention of these approximation schemes. The Homogeneous Electron Fluid (HEF) is the basic model on which condensed matter physics is built. Density Functional Theory (DFT) is an attempt to write down phenomenological ansatze for the correlation functions of this field theory, in order to understand real materials. From the earliest days of study of the HEF it was realized that it had both a fluid and a crystalline phase. More recentlyks colloidal phases have been proposed, and a persuasive mean field argument for the existence of these phases has been given in two dimensions.
In a series of papers tbbz1; tbbz2; tbbz3, one of the authors has tried to explore these colloidal phases more systematically. In banks2020systematic, he proposed that one could do so by searching for crystalline solutions of the truncated large $N$ SD equations. These would provide a more systematic approach to the Wigner crystal phase than the complicated semi-classical solution of the many body Schrodinger equation and allow an extrapolation of the solution into the region where the crystal became meta-stable. As argued in tbbz1; tbbz2; tbbz3 this is the regime where one expects gel like colloids to exist. One should also be able to follow the system through its metal-insulator transition and check whether the Kivelson-Spivak prediction of crystalline impurities with negative surface tension is valid at large enough $N$. Finally, the disappearance of the crystalline solution at a finite value of the density would be a neat explanation of the mysterious phase transition unveiled in haule; japan.
@@
Table A1: Table of numerically evaluated 1 and 2-pt correlation functions for $M=9$ and $N=40,100,$ and $1000$, along with the right hand side expressions of eqs. (4) and (5) with these values plugged in. Index $k$ corresponds to $\langle x(k)\rangle$ in the first three columns and eq. (4) with $i=k$ in the last three columns. Similarly, index klkl corresponds to $\langle x(k)x(l)\rangle$ in the first three columns and eq. (4) with $i=k$ and $t=l$ in the last three columns. It is easy to that the 1-pt functions and the equation results, on average, grow closer to 0 for larger $N$.